OpenStudy (anonymous):

what is 15 centimetresquare to metre square

OpenStudy (anonymous):

If you cannot quite see it yet, try doing this:\[\ 1cm^2 =1cm*cm\]and \[\ 1cm=10^{-2} m\]so \[\ 1cm^2=1*10^{-2}m*10^{-2}m =1*10^{-4}m^2 \] now it is easy to see that \[\ cm^2=10^{-4}m^2\]

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):